OpenStudy (anonymous):

A man is currently 20yrs older than his oldest son. His oldest son is currently two yrs older than his youngest son. Three years from the current time, the product of the ages of the brothers will be ten more than eleven times the age of the father. What is the current age of the man's youngest son?

hero (hero):

f = o + 20 o = y + 2 (o+3)(y+3) = 11f + 10 y = ?

OpenStudy (anonymous):

f=0 or O?

OpenStudy (anonymous):

are these all one equation or different equations? I am trying to figure it out by substitute lol not working

hero (hero):

Wait...

OpenStudy (anonymous):

okay

hero (hero):

Actually, it is not solvable.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

it's the same kind of problem I was running into when trying to substitute

hero (hero):

Well, if it's solveable, you'll have to use the quadratic formula to solve it

OpenStudy (anonymous):

with all your equal signs I can't tell if it's continous or just different equations not broken up

hero (hero):

y^2 - 6y -246 = 0

OpenStudy (anonymous):

that looks better

hero (hero):

There you go. Plug that into the quad form

hero (hero):

Warning: The values will not be "even"

OpenStudy (anonymous):

it can be both 3+ squareroot of 1020/2 or 3-squareroot of 1020-2

hero (hero):

If that is correct, you will need to round it

hero (hero):

to decimal

OpenStudy (anonymous):

ugh

hero (hero):

lol

OpenStudy (anonymous):

Thanks

hero (hero):

What did you get?

OpenStudy (anonymous):

about 15.97